# Lips Are Lying: Spotting the Temporal Inconsistency between Audio and Visual in Lip-Syncing DeepFakes

### v1

@@ Abstract
In recent years, DeepFake technology has achieved unprecedented success in high-quality video synthesis, whereas these methods also pose potential and severe security threats to humanity. DeepFake can be bifurcated into entertainment applications like face swapping and illicit uses such as lip-syncing fraud. However, lip-forgery videos, which neither change identity nor have discernible visual artifacts, present a formidable challenge to existing DeepFake detection methods. Our preliminary experiments have shown that the effectiveness of the existing methods often drastically decreases or even fails when tackling lip-syncing videos.
In this paper, for the first time, we propose a novel approach dedicated to lip-forgery identification that exploits the inconsistency between lip movements and audio signals. We also mimic human natural cognition by capturing subtle biological links between lips and head regions to boost accuracy. To better illustrate the effectiveness and advances of our proposed method, we curate a high-quality LipSync dataset by employing the SOTA lip generator. We hope this high-quality and diverse dataset could be well served the further research on this challenging and interesting field. Experimental results show that our approach gives an average accuracy of more than 95.3% in spotting lip-syncing videos, significantly outperforming the baselines. Extensive experiments demonstrate the capability to tackle deepfakes and the robustness in surviving diverse input transformations. Our method achieves an accuracy of up to 90.2% in real-world scenarios (e.g., WeChat video call) and shows its powerful capabilities in real scenario deployment. To facilitate the progress of this research community, we release all resources at https://github.com/AaronComo/LipFD.
Figure 1: A visualization comparison between common deepfakes involves facial attribute editing and our studied lip-syncing deepfakes (LipSync). The former exhibits a substantial forgery area and identity manipulation, such as face swapping, whereas the latter, lip-syncing deepfakes, relies on the synchronization of the minor lip region and given audio, without any alterations to the subject’s identity. As illustrated in the comparison above, discerning the authenticity of an image sequence becomes arduous in the absence of accompanying labels.
footnotetext: ${}^{\dagger}$qual contribution.footnotetext: $*{}^{*}$orresponding author. E-mail: wangrun@whu.edu.cn
@@ 1 Introduction
DeepFake refers to an AI-based technology for synthesizing fake media data Dolhansky et al. (2020). The recent advancements in generative models, particularly the emergence of several GAN architectures Karras et al. (2019); Liu et al. (2019); Choi et al. (2018) and the diffusion probabilistic models Ho et al. (2020), have enhanced the realism and quality of forged videos that can easily deceive humans. The prevalence of DeepFake poses potential security risks, e.g., political elections and identity verification, sparking public concerns.
DeepFake can be bifurcated into entertainment applications and illicit uses Juefei-Xu et al. (2022). As illustrated in fig. 1, the popular DeepFake aims to bring fun to users by swapping face identity information to synthesize new content, such as gender swapping and age regression. Unfortunately, the severe DeepFake is utilized for illicit crimes, including manipulating political propaganda and fabricating pornographic content. The case is particularly alarming in LipSync fraud, where the audio drives the mouth movements on reconstructed video frames. These DeepFakes are generally exploited by malicious actors in real-world scenarios, such as the widely disseminated fabricated videos of Barack Obama saying things he never said on YouTube Suwajanakorn et al. (2017), posing significant security threats. The escalating issue of real-time forgery necessitates an effective detector to identify videos generated through LipSync.
@@
Existing studies on DeepFake detection can be classified into unimodal-based and multi-modal-based methods, where the former relies on visual discrepancies arising from identity tampering to detect Lutz and Bassett (2021); Zhao et al. (2020). However, unimodal detectors become less reliable when the forged videos are perturbed for targeted removal of LipSync artifacts. In recent years, several multi-modal-based methods have emerged Cozzolino et al. (2023); Yang et al. (2023); Hashmi et al. (2022), including audio-visual fusion and audio-visual inconsistency. Fusion strategies may confound the learning of singular modality features and the performance post-fusion is not necessarily enhanced Khalid et al. (2021). Muppalla et al. (2023) suggested training detectors to learn the inconsistencies between video frames and audio. However, as the arms race between DeepFake creation and detection intensifies, these inconsistencies are gradually reduced, making coarse-grained audio-visual alignment strategies less effective against advanced LipSync methods.
Lip movements are discrete, while the audio spectrum is continuous, resulting in inherent inconsistencies in LipSync videos. As illustrated in the fig. 3, we observe a temporal correlation between the energy variations in audio spectrum and lip movements. To the best of our knowledge, existing works naively align single-frame images with long-range audio clips, thus neglecting the temporal inconsistencies of audio-visual features Agarwal et al. (2020); Mittal et al. (2020). Our experimental evidence also indicates a marked decline in the efficacy of existing methods when confronting LipSync forged videos. Moreover, Kotsia et al. (2008) demonstrated the mouth region’s significance in facial appearance, surpassing even the eyes, due to its biological connections with other head regions. Humans naturally leverage the cues of local regions and head postures to discern facial semantics. While DeepFake technology has made strides in replicating overall facial dynamics, it often falls short of accurately simulating these subtle yet crucial biological interactions. Hence, we choose to exploit the biologically intrinsic correlation between lip movements and head postures as auxiliary information to detect deepfakes. This approach not only mimics the natural cognitive processes of humans but also capitalizes on the existing limitations of deepfakes.
@@
To evaluate the effectiveness and generalization of our approach in detecting lip-syncing deepfakes, we utilize the state-of-the-art LipSync methods to generate massive high-quality lip forgery video dataset based on Lip Reading Sentences 2 (LRS2) Afouras et al. (2018), Face Forensics++ (FF++) Rossler et al. (2019), Deepfake Detection Challenge Dataset (DFDC) Dolhansky et al. (2020). Experimental results show that our approach outperforms prior works by a notable margin, with an accuracy up to 96.93% for four types of lip forgery videos. Rigorous ablations of our design choices and comparisons with other detection methods demonstrate the superiority of our approach. To summarize, our main contributions are summarized as follows:
@@
  • In this work, we unveil a key insight that exploits the discrepancies between lip movements and audio signals for fine-grained forgery detection. Our approach introduces a dual-headed detection architecture to enhance detection capabilities.
  • We construct the first large scale audio-visual LipSync dataset with nearly one hundred thousand samples, and conducted comprehensive experiments on it alongside other DeepFake datasets. Our method demonstrated high efficacy and robustness, achieving around 94% average accuracy in LipSync detection, and up to 90.18% in real-world scenarios.
@@ 2 Related Work
Lip-syncing Generation. Lip-syncing facial manipulation, which forges a speaker’s lip movements to match a given audio, is among the most threatening DeepFake applications due to its subtlety and difficulty to detect, typically falsifying the speaker’s conveyed information. Zhou et al. (2020) disentangled the content and speaker information in the audio signal, allowing attackers to generate a forged video using just a single image and an audio segment. Still, it is weak in representing bilabial and fricative sounds due to the omission of short phoneme representations. Prajwal et al. (2020) introduced a well-trained discriminator and a temporal consistency checker to address the loss of short-duration phoneme, enhancing the authenticity of generated videos. However, it exhibits weak temporal coordination in the lip movement of talking heads. Wang et al. (2023) further focuses on the content of lip movements, making the forgeries challenging for both human eyes and machines to recognize.
@@
DeepFake Detection. The existing DeepFake detectors employ single-modal-based or multi-modal-based approaches to detect subtle differences between real and fake samples. Earlier single-modal detectors aspired to employ neural networks to automatically extract discriminative information Wodajo and Atnafu (2021); Zhao et al. (2021), but they failed to detect unseen samples due to overfitting. To address this issue, some studies shift focus to frequency domain features Liu et al. (2021) or subtle forgery artifacts in more generalized datasets Chen et al. (2021); Shiohara and Yamasaki (2022). Another line is to guide the network to focus on discriminative locations, such as automatically guiding the detector’s positional attention through a double-stream network Shuai et al. (2023), or manually cropping the lip region to extract artifacts formed by the inconsistent lip movements Haliassos et al. (2021). Although these works have achieved considerable performance on afore datasets, they are not sensitive when faced with advanced lip-syncing generators due to the absence of synchronized audio features. In the multi-modal-based detectors, noticing that the coordination of audio-visual modalities is an inherently challenging issue in any SOTA generator, Chugh et al. (2020) quantifies the disparity between audio and visual as the criterion for classification, but focusing too much on the background information in the video led to failure. In this context, Haliassos et al. (2022) intentionally extracts talking head movements and establishes a correlation with audio for discrimination. These methods performed well in addressing audio-visual forgery, but are susceptible to the influence of noise or compression.
Figure 2: Real relation between lip movements and spectrogram. When the woman starts talking, the middle and high frequencies in the spectrum are lighted. As time passes, the energy gradually fades and shifts from the middle frequencies to lower frequencies.
Figure 2: Real relation between lip movements and spectrogram. When the woman starts talking, the middle and high frequencies in the spectrum are lighted. As time passes, the energy gradually fades and shifts from the middle frequencies to lower frequencies.
Figure 3: Lip-syncing fake videos. The first two frames show a highlighted high-frequency spectrum, which is contradictory to the man not speaking. In the third frame, an unexpected opening of lip appears at the darkest part in the spectrum. On one hand, the mouth cannot change so drastically within a single frame. On the other hand, this lip’s shape contradicts the spectrum information.
@@ 3 LipSync Forgery Dataset
To the best of our knowledge, the majority of public DeepFake datasets consist solely of video or image sources, with no specialized dataset specifically dedicated to lip forgery detection available. To fill this gap, we construct a high-quality audio-visual dataset, AVLips, which contains 100,000 audio-visual samples generated by several state-of-the-art LipSync methods. The workflow is demonstrated in fig. 4.
High quality. We employed a combination of static “MakeItTalk” Zhou et al. (2020) and dynamic “Wav2Lip” Prajwal et al. (2020), “TalkLip” Wang et al. (2023) generation methods to simulate realistic lip movements. These generation methods are widely recognized as high-quality work in the academic community, capable of generating high-resolution videos while ensuring accurate lip movements. We applied a noise reduction algorithm to all audio samples before synthesis to reduce irrelevant background noise, ensuring the models can focus on speech content.
Diversity. Our dataset encompasses a wide range of scenarios, covering not only well-known public datasets but also real-world WeChat video calls. Our aim is for this collection to act as a catalyst for advancing real-time forgery detection. To better simulate the nuances of real-world conditions, we have employed six perturbation techniques — saturation, contrast, compression, Gaussian noise, Gaussian blur, and pixelation — at various degrees, thus ensuring the dataset’s realism and practical relevance.
@@ 4 Method
In reality, the movements of a speaker’s lip and head are closely intertwined with the spoken content, forming a natural and coherent unity. These physical movements naturally align with the timing and context of the speech. However, the LipSync method, which solely relies on audio signals to generate lip movements frame by frame, only focuses on the precise alignment between lip shapes and speech at any given moment. It overlooks the broader temporal context and the overall coherence of lip and head movements during speech. Consequently, the generated audio-visual outputs often exhibit inherent inconsistencies regarding temporal synchronization. These inconsistencies serve as valuable clues and insights for our detection efforts, highlighting the disparity between natural lip movements and artificially generated ones. Fig. 3 and 3 vividly exhibit the temporal features among those two classes.
Extracting temporal inconsistencies between audio and video presents notable challenges due to the utilization of features from multiple modalities. To tackle this, we developed a dual-headed detection architecture presented in fig. 5. (1) The Global Feature encoder is dedicated to encoding temporal features, capturing the overarching correlation between audio and lip movements. (2) The Global-Region encoder aims to detect subtle visual forgery traces within regions of varying scales and integrate them with global features. (3) Moreover, we introduced an innovative Region Awareness module that dynamically adjusts the model’s attention across different scales. We will demonstrate in section 6.2 that this module stands as a cornerstone, harnessing features from regions of diverse sizes, thus empowering our model to effectively capture both the prominent changes in DeepFake and the subtle adjustments in LipSync.
@@ 4.1 Global Feature Encoding
Based on the findings mentioned before, we need to extract features in the temporal domain. Inspired by the translation task in natural language processing, where transformers detect long-distance vocabulary correlations, we regard the inherent correlation between lip movements and spectral information as analogous to the relationship between “vocabulary” in a “sentence” sequence. To capture and encode this correlation, we employ a transformer model.
To effectively carry out its task, the encoder necessitates extraordinary representational capacity, which can be attained through exposure to a vast number of images Ojha et al. (2023). This capacity enables the encoder to accurately allocate attention to the relevant regions of interest. To satisfy this requirement, we choose a variant of vision transformer ViT:L/14 Dosovitskiy et al. (2020), pre-trained on CLIP Radford et al. (2021). In our experiments, we use the final layer of CLIP: ViT-L/14’s visual encoder for image embedding.
Formulation. We denote the convolutional layer as $Conv$, which convolves images down to $224\times 224$. We first crop source image $I$ into 3 series as $\{c_{h}^{N},c_{f}^{N},c_{l}^{N}\}_{i},\ i\in\{0,...,T-1\}$, where $N$ equals to batch size and $T$ notes the window size. Image $I$ will be embedded into $F_{G}$ as global feature:
$F_{G}=ViT(Conv(I,\theta_{Conv}))$
$\small{\{c_{h}^{N},c_{f}^{N},c_{l}^{N}\}_{i}=Crop(I,\{1.0,0.65,0.45\}),\ i\in\{0,1,2\}}$ (2)
where $\theta_{Conv}$ is the parameters of $Conv$. The encoder is constrained by $\mathcal{L}_{RA}$ that is to be further described in the following section.

### v2

@@ Abstract
In recent years, DeepFake technology has achieved unprecedented success in high-quality video synthesis, but these methods also pose potential and severe security threats to humanity. DeepFake can be bifurcated into entertainment applications like face swapping and illicit uses such as lip-syncing fraud. However, lip-forgery videos, which neither change identity nor have discernible visual artifacts, present a formidable challenge to existing DeepFake detection methods. Our preliminary experiments have shown that the effectiveness of the existing methods often drastically decrease or even fail when tackling lip-syncing videos. In this paper, for the first time, we propose a novel approach dedicated to lip-forgery identification that exploits the inconsistency between lip movements and audio signals. We also mimic human natural cognition by capturing subtle biological links between lips and head regions to boost accuracy. To better illustrate the effectiveness and advances of our proposed method, we create a high-quality LipSync dataset, AVLips, by employing the state-of-the-art lip generators. We hope this high-quality and diverse dataset could be well served the further research on this challenging and interesting field. Experimental results show that our approach gives an average accuracy of more than 95.3% in spotting lip-syncing videos, significantly outperforming the baselines. Extensive experiments demonstrate the capability to tackle deepfakes and the robustness in surviving diverse input transformations. Our method achieves an accuracy of up to 90.2% in real-world scenarios (e.g., WeChat video call) and shows its powerful capabilities in real scenario deployment. To facilitate the progress of this research community, we release all resources at https://github.com/AaronComo/LipFD.
Figure 1: A visualization comparison between common deepfakes and our studied lip-syncing deepfakes (LipSync). The former exhibits a substantial forgery area and identity manipulation, such as face or gender swapping, whereas the latter, relies on the synchronization of the minor lip region and given audio, without any alterations to the subject’s identity. As illustrated in the comparison above, discerning the authenticity of an image sequence becomes arduous in the absence of labels.
@@ 1 Introduction
DeepFake refers to an AI-based technology for synthesizing fake media data dfdc . The recent advancements in generative models, particularly the emergence of several GAN architectures karras2019style ; liu2019stgan ; choi2018stargan and the diffusion probabilistic models ho2020denoising , have enhanced the realism and quality of forged videos that can easily deceive humans. The prevalence of DeepFake poses potential security risks, e.g., political elections and identity verification, sparking public concerns wang2022anti .
DeepFake can be bifurcated into entertainment applications and illicit uses juefei2022countering . As illustrated in Fig. 1, the popular DeepFake aims to bring fun to users by swapping faces to synthesize new content, such as gender swapping and age regression. Unfortunately, the severe DeepFake is utilized for illicit crimes, including manipulating political propaganda and fabricating pornographic content. The case is particularly alarming in LipSync fraud, where the audio drives the mouth movements on reconstructed video frames. These DeepFakes are generally exploited by malicious actors in real-world scenarios, such as the widely disseminated fabricated videos of Barack Obama saying things he never said on YouTube suwajanakorn2017synthesizing , posing significant security threats. The escalating issue of real-time forgery necessitates an effective detector to identify videos generated through LipSync.
@@
Existing studies on DeepFake detection can be classified into unimodal-based and multi-modal-based methods, where the former relies on visual discrepancies arising from identity tampering to detect lutz2021deepfake ; zhao2020capturing ; liang2024let ; Juefei-Xu2022-ug . However, unimodal detectors become less reliable when the forged videos are perturbed for targeted removal of LipSync artifacts. In recent years, several multi-modal-based methods have emerged cozzolino2023audio ; yang2023avoid ; hashmi2022multimodal , including audio-visual fusion and audio-visual inconsistency. Fusion strategies may confound the learning of singular modality features and the performance post-fusion is not necessarily enhanced khalid2021fakeavceleb . muppalla2023integrating suggested training detectors to learn the inconsistencies between video frames and audio. However, as the arms race between DeepFake creation and detection intensifies, these inconsistencies are gradually reduced, making coarse-grained audio-visual alignment strategies less effective against advanced LipSync methods.
Lip movements are discrete, while the audio spectrum is continuous, resulting in inherent inconsistencies in LipSync videos. As illustrated in Fig. 2, we observe a temporal correlation between the energy variations in spectrum and lip movements. To the best of our knowledge, existing works naively align single-frame images with long-range audio clips, thus neglecting the temporal inconsistencies of audio-visual features 9151013 ; mittal2020emotions . Our experimental evidence also indicates a marked decline in the efficacy of existing methods when confronting LipSync forged videos. Moreover, kotsia2008analysis demonstrated the mouth region’s significance in facial appearance, surpassing even the eyes, due to its biological connections with other head regions. Humans naturally leverage the cues of local regions and head postures to discern facial semantics. While DeepFake technology has made strides in replicating overall facial dynamics, it often falls short of accurately simulating these subtle yet crucial biological interactions. Hence, we choose to exploit the biologically intrinsic correlation between lip movements and head postures as auxiliary information to detect deepfakes. This approach not only mimics the natural cognitive processes of humans but also capitalizes on the existing limitations of deepfakes.
@@
To evaluate the effectiveness and generalization of our approach in detecting lip-syncing deepfakes, we utilize the state-of-the-art LipSync methods to generate massive high-quality lip forgery video dataset based on Lip Reading Sentences 3 (LRS3) lrs3 , Face Forensics++ (FF++) ff++ , Deepfake Detection Challenge Dataset (DFDC) dfdc . Experimental results show that our approach outperforms prior works by a notable margin, with an accuracy up to 96.93% for four types of lip forgery videos. Rigorous ablations of our design choices and comparisons with other detection methods demonstrate the superiority of our approach. Our main contributions can be summarized as follows:
@@
  • In this work, we unveil a key insight that exploits the discrepancies between lip movements and audio signals for fine-grained forgery detection. Our approach introduces a dual-headed model architecture to enhance detection capabilities.
  • We construct the first large scale audio-visual LipSync dataset with up to 340,000 samples, and conducted comprehensive experiments on it alongside other DeepFake datasets. Our method demonstrated high efficacy and robustness, achieving around 95% average accuracy in LipSync detection, and up to 90.18% in real-world scenarios.
Figure 2: (a) shows the correlation between lip movements and corresponding spectrogram in genuine pattern. When the woman starts talking, the middle and high frequencies in the spectrum are lighted. Over time, the energy gradually fades and shifts from middle to lower frequencies. (b) the first two frames show a highlighted high-frequency spectrum, contradicting the man not speaking. In the third frame, an unexpected lip opening appears at the darkest part of the spectrum. The mouth cannot change so drastically within a single frame, and this lip shape contradicts the spectrum information.
@@ 2 Related Work
Lip-syncing Generation. Lip-syncing facial manipulation, which forges a speaker’s lip movements to match a given audio, is among the most threatening DeepFake applications due to its subtlety and difficulty to detect, typically falsifying the speaker’s conveyed information. makeittlak disentangled the content and speaker information in the audio signal, allowing attackers to generate a forged video using just a single image and an audio segment. Still, it is weak in representing bilabial and fricative sounds due to the omission of short phoneme representations. Wav2Lip introduced a well-trained discriminator and a temporal consistency checker to address the loss of short-duration phoneme, enhancing the authenticity of generated videos. However, it exhibits weak temporal coordination in the lip movement of talking heads. talklip further focuses on the content of lip movements, making the forgeries challenging for both human eyes and machines to recognize.
@@
DeepFake Detection. The existing DeepFake detectors employ single-modal-based or multi-modal-based approaches to detect subtle differences between real and fake samples. Earlier single-modal detectors aspired to employ neural networks to automatically extract discriminative information wodajo2021deepfake ; zhao2021multi , but they failed to detect unseen samples due to overfitting. To address this issue, some studies shift focus to frequency domain features liu2021spatial ; Wu_Ma_Wang_Zhang_Liang_Li_Lin_Fang_Wang_2024 or subtle forgery artifacts in more generalized datasets chen2021local ; shiohara2022detecting . Another line is to guide the network to focus on discriminative locations, such as automatically guiding the detector’s positional attention through a double-stream network shuai2023locate , or manually cropping the lip region to extract artifacts formed by the inconsistent lip movements lipforensic . Although these works have achieved considerable performance on afore datasets, they are not sensitive when faced with advanced lip-syncing generators due to the absence of synchronized audio features. In the multi-modal-based detectors, noticing that the coordination of audio-visual modalities is an inherently challenging issue in any SOTA generator, chugh2020not quantifies the disparity between audio and visual as the criterion for classification, but focusing too much on the background information in the video led to failure. In this context, haliassos2022leveraging intentionally extracts talking head movements and establishes a correlation with audio for discrimination. These methods performed well in addressing audio-visual forgery, but are susceptible to the influence of noise or compression.
@@ 3 LipSync Forgery Dataset
To the best of our knowledge, the majority of public DeepFake datasets consist solely of videos or images, with no specialized one specifically dedicated to LipSync detection available. To fill this gap, we construct a high-quality Audio-Visual Lip-syncing Dataset, AVLips, which contains up to 340,000 audio-visual samples generated by several SOTA LipSync methods. The workflow is demonstrated in Fig. 3.
High quality. We employed a combination of static ‘MakeItTalk’ makeittlak and dynamic ‘Wav2Lip’ Wav2Lip , ‘TalkLip’ talklip , ‘SadTalker’ zhang2023sadtalker generation methods to simulate realistic lip movements. These methods are widely recognized as high-quality work, capable of generating high-resolution videos while ensuring accurate lip movements. We applied a noise reduction algorithm to all audio samples before synthesis to reduce irrelevant background noise, ensuring the models can focus on speech content.
Diversity. Our dataset encompasses a wide range of scenarios, covering not only well-known public datasets but also real-world data. Our aim is for this collection to act as a catalyst for advancing real-time forgery detection. To better simulate the nuances of real-world conditions, we have employed six perturbation techniques — saturation, contrast, compression, Gaussian noise, Gaussian blur, and pixelation — at various degrees, thus ensuring the dataset’s realism and practical relevance.
@@ 4 Method
In reality, the movements of a speaker’s lip and head are closely intertwined with the spoken content, forming a natural and coherent unity. These physical movements naturally align with the timing and context of the speech. However, LipSync method, which solely relies on audio signals to generate lip movements frame by frame, only focuses on the precise alignment between lip shapes and speech at any given moment. It overlooks the broader temporal context and the overall coherence of lip and head movements during speech. Consequently, the generated outputs often exhibit inherent inconsistencies regarding temporal synchronization. These inconsistencies serve as valuable clues and insights for our detection efforts, highlighting the disparity between natural lip movements and artificially generated ones. Fig. 2 vividly exhibit the temporal features among those two classes.
Extracting temporal inconsistencies between audio and video presents notable challenges due to the utilization of features from multiple modalities. To tackle this, we developed a dual-headed detection architecture presented in Fig. 4. (1) The Global Feature encoder is dedicated to encoding temporal features, capturing the overarching correlation between audio and lip movements. (2) The Global-Region encoder aims to detect subtle visual forgery traces within regions of varying scales and integrate them with global features. (3) Moreover, we introduced an innovative Region Awareness module that dynamically adjusts the model’s attention across different scales. We will demonstrate in Sec. 8 that this module stands as a cornerstone, harnessing features from regions of diverse sizes, thus empowering our model to effectively capture both the prominent changes in DeepFake and the subtle adjustments in LipSync.
@@ 4.1 Global Feature Encoding
Based on the findings mentioned before, we need to extract features in the temporal domain. Inspired by the translation task in natural language processing, where transformers detect long-distance vocabulary correlations, we regard the inherent correlation between lip movements and spectral information as analogous to the relationship between ‘vocabulary’ in a ‘sentence’ sequence. To capture and encode this correlation, we employ a transformer model.
To effectively carry out its task, the encoder necessitates extraordinary representational capacity, which can be attained through exposure to a vast number of images UniversalFakeDetect . This capacity enables the encoder to accurately allocate attention to the relevant regions of interest. To satisfy this requirement, we choose a variant of vision transformer ViT:L/14 vit , pre-trained on CLIP clip . In our experiments, we use the final layer of CLIP: ViT-L/14’s visual encoder for image embedding.
Formulation. We denote the convolutional layer as $Conv$, which convolves images down to $224\times 224$. We first crop source image $I$ into 3 series as $\{c_{h}^{N},c_{f}^{N},c_{l}^{N}\}_{i},\ i\in\{0,...,T-1\}$, where $N$ equals to batch size $T$ notes the window size, $c_{l}$ is the lip region subject to modifications by LipSync, $c_{f}$ represents face area focused on by DeepFake, and $c_{h}$ encompasses the overall zone containing head posture and background information. The three-tiered cropping strategy emulate the human visual focus on key facial areas, spotlighting the lip and overall facial structure. Image $I$ will be embedded into $F_{G}$ as global feature:
$F_{G}=ViT(Conv(I))$ (1)
$\{c_{h}^{N},c_{f}^{N},c_{l}^{N}\}_{i}=Crop(I,\{1.0,0.65,0.45\}),\ i\in\{0,1,2\}$ (2)
The encoder is constrained by $\mathcal{L}_{RA}$ that is to be further described in the following section.